# Logic-in-Frames: Dynamic Keyframe Search via Visual Semantic-Logical Verification for Long Video Understanding

### v1

@@ 3.1 Benchmark Datasets
The proposed VSLS is systematically evaluated across four benchmark datasets: a) LongVideoBench [38] for assessing long-context video-language comprehension capabilities; b) Video-MME [11] as the first comprehensive benchmark for multimodal video analytics; c) Haystack-LVBench, extended from LongVideoBench with human-annotated frame index answers; and d) Haystack-Ego4D, derived from Ego4D with similar annotations.
While LongVideoBench and Video-MME measure performance enhancement in QA accuracy, Haystack-Ego4D and Haystack-LVBench quantitatively evaluate keyframe selection accuracy through recall and precision metrics. Further details of these datasets, including descriptions, sizes, and usages are provided in Appendix C.
@@ 3.2 Evaluation Metrics
LongVideoBench Video-MME Model and Size Frame Video Length Model and Size Frame Video Length Long Medium Short Long Medium Short 900-3600s 180-600s 15-60s 30-60min 4-15min 0-2min GPT-4o [14] 8 47.1 49.4 67.3 GPT-4o 8 55.2 60.2 69.6 GPT-4o + T* 8 49.1 56.2 68.0 GPT-4o + T* 8 55.2 61.2 68.9 GPT-4o + VSLS (ours) 8 51.2 58.9 74 GPT-4o + VSLS (ours) 8 56.9 60.7 68.2 InternVL 2.5-78B [5] 8 55.7 57.3 74.0 InternVL 2.5-78B 8 52.6 55.5 55.9 InternVL 2.5-78B + VSLS (ours) 8 58.0 61.5 74.0 InternVL 2.5-78B + VSLS (ours) 8 57.7 57.5 59.0 GPT-4o 32 53.8 56.5 74.0 GPT-4o 32 55.2 61.0 71.4 GPT-4o + T* 32 55.3 58.8 72.0 GPT-4o + T* 32 55.2 61.6 72.6 GPT-4o + VSLS (ours) 32 54.2 60.0 76.0 GPT-4o + VSLS (ours) 32 55.2 61.9 71.9 LLaVA-OneVision-QWen2-78B-ov 32 59.3 63.9 77.4 LLaVA-OneVision-78B 32 60.0 62.2 66.3 PLLaVA-34B 32 49.1 50.8 66.8 VideoLLaMA 2 32 57.6 59.9 62.4 LLaVA-Video-78B-Qwen2 128 59.3 63.9 77.4 Oryx-1.5 128 59.3 65.3 67.3 mPLUG-Owl3-7B 128 53.9 58.8 73.7 Aria-8x3.5B 256 58.8 67.0 67.6 GPT-4o (0513) 256 61.6 66.7 76.8 Gemini-1.5-Pro (0615) 1/0.5 fps 67.4 74.3 75.0
Table 2: Downstream task evaluation results on two benchmarks. All the accuracy(%) in black is based on our replication results. We also cite the reported accuracy of the SOTA models in gray (note that their settings may differ from ours, and these results are not guaranteed to be reproducible) and list their number of frames used for QA inference, which we include for full transparency and openness.
@@ 3.2.1 Evaluation Metrics for Search Utility
Our assessment framework emphasizes both the effectiveness and efficiency of search functions. In terms of search effectiveness, we use three metrics to compare model-predicted keyframes with human annotations, considering both individual frames and entire sets. This approach tackles the issue that multiple valid keyframe sets might exist for a single query. For frame-to-frame comparisons, we evaluate the alignment between a model-predicted frame $f_{\text{pt}}$ and a human-annotated frame $f_{\text{gt}}$ from two perspectives:
@@
where $T_{\text{pt}}$ represents the set of predicted timestamps and $T_{\text{gt}}$ represents the set of ground truth timestamps. In this formula, $|N_{\text{gt}}|$ denotes the total number of ground truth frames, $t_{\text{gt}}^{i}$ is the timestamp of the $i$-th ground truth frame, $t_{\text{pt}}^{j}$ is the timestamp of the $j$-th predicted frame, $\delta$ is the temporal similarity threshold defining the maximum acceptable time difference, and $\mathbb{I}[\cdot]$ is the indicator function that returns 1 when the condition is satisfied and 0 otherwise.
Visual Similarity uses the Structural Similarity Index Measure (SSIM) [1] to evaluate the visual resemblance between $f_{\text{pt}}$ and $f_{\text{gt}}$ based on structural details, luminance, and contrast. For set-to-set comparisons, the main challenge is defining the similarity between two sets. We use Precision $P$ and Recall $R$ as complementary metrics. Precision checks if each model-selected frame matches at least one reference frame, while Recall ensures that reference frames are adequately represented in the model’s output. Given the ground truth keyframe set $F_{\text{gt}}=\{f^{j}_{\text{gt}}\}^{n}_{j=1}$ and the predicted keyframe set $F_{\text{pt}}=\{f^{j}_{\text{pt}}\}^{m}_{i=1}$, we formally define the multimodal retrieval quality metrics as follows:
$P(F_{\text{pt}},F_{\text{gt}})$ $=\frac{1}{|F_{\text{pt}}|}\sum_{f^{i}_{\text{pt}}\in F_{\text{pt}}}\max_{f^{j}_{\text{gt}}\in F_{\text{gt}}}\phi(f^{i}_{\text{pt}},f^{j}_{\text{gt}});$ (12a)
@@ 3.2.2 Evaluation Metrics for Search efficiency
Existing studies [10, 21, 31, 33, 36] have mainly concentrated on optimizing task-specific performance metrics while neglecting computational efficiency in temporal search operations. To systematically analyze this dimension, our evaluation framework incorporates two criteria: 1) FLOPs representing arithmetic operation complexity, and 2) Latency recording real-world execution duration.
@@ 3.4 Visual Semantic Logic Search Performance
These results empirically prove the foundational hypothesis that extracting semantic and logical cues from textual queries facilitates precise detection of query-relevant video frames. The marked improvements in visual similarity and temporal coverage demonstrate that VSLS successfully captures corresponding keyframes while maintaining temporal coherence through the visual-logical alignment.
Method Frame LongVideoBench Precision $\uparrow$ Recall $\uparrow$ Time $\uparrow$ Static Frame Sampling Method Uniform [38] 8 56.0 72.0 6.3 Uniform 8 60.7 80.4 4.7 Uniform 32 58.7 81.6 24.9 Uniform 32 60.2 85.0 8.1 Dense Retrieval Method VideoAgent [10] 10.1 58.8 73.2 8.5 Retrieval-based [39] 8 63.1 65.5 6.3 Retrieval-based 32 59.9 80.8 21.8 Temporal Searching Method T* [39] 8 58.4 72.7 7.1 T* 8 75.3 88.2 26.2 VSLS (ours) 8 75.6 88.6 26.3 T* 32 58.3 83.2 28.2 T* 32 74.0 90.3 36.5 VSLS (ours) 32 74.5 92.5 41.4
Table 3: Results of searching utility on LongVideoBench. 8-frame setting bests are underlined, and 32-frame setting bests are in bold. Gray font indicates results from the original paper.
@@ 3.5 Downstream Video QA Performance
To further demonstrate the advantages of our VSLS, we conduct downstream video QA performance experiments on LongVideoBench and Video-MME datasets. As shown in Table 2, we categorize the video length into three groups, namely Long, Medium, and Short, ranging from 15 to 3600s, and 0 to 60 mins respectively. Our method consistently achieves the highest accuracy in the long video category across different numbers of frames and downstream QA models. Compared to our baseline method T*, we observe significant performance improvements by incorporating our selected visual semantic logic relation, which is shown in Figure 1. These results validate that our approach can capture and utilize visual logical relations in long video content, leading to superior video QA performance.
4 Analysis
4.1 Time Complexity
The proposed framework can be divided into two stages. In the initial phase, VLMs such as LlaVA-7B and GPT-4o, are employed to extract a semantic set $\mathcal{S}$ from a video $V$ containing $n$ frames. This semantic set $\mathcal{S}$ comprises three semantic components: target objects, cue objects, and their relational information. Notably, the size of $\mathcal{S}$ is constrained by our prompt design. Following this, the subsequent phase focuses on key-frame identification and operates as a heuristic search algorithm. Specifically, this process iteratively selects $k$ candidates from the $n$ frames through a scoring mechanism guided by an evaluation function $h(\cdot,\mathcal{S})$. Furthermore, the score distribution $\text{scores}[n]$ is dynamically updated based on the YOLO-World detector’s output, ensuring continuous refinement of the selection process.
Our analysis will focus on the number of YOLO-World detections required during the search process, which is the primary computational bottleneck of our framework. This is because each detection involves running a deep neural network, making it the most resource-intensive step. Reducing the number of detections directly improves efficiency without compromising accuracy. In each iteration, the YOLO-World detector is leveraged to match the selected $k$ frames with the objects and relations in $\mathcal{S}$, resulting in $k$ YOLO-World detections. The search terminates when all objects are detected and their relations are satisfied or when the number of iterations reaches the predefined search budget, defined as $\mathrm{min}(1000,0.1\times V_{t})$, where $V_{t}$ denotes the duration of the video $V$ in seconds. In the worst cases, for a video longer than 10,000 frames where each iteration fails to identify any object and relation, the number of iterations will be clamped at 1,000. Under ideal conditions, the evaluation function $h(\cdot,\mathcal{S})$ consistently identifies all target frames and leads to the optimal score distribution; the algorithm is similar to selecting top-$k$ candidates from n frames within $\mathcal{O}(|\mathcal{S}|\log n)$ iterations [39]. Therefore, the average time complexity of our algorithm is $\mathcal{O}(|\mathcal{S}|k\log n)$ YOLO-World inferences.
Experimental results also demonstrate that integrating relational information into the search algorithm incurs negligible computational overhead compared to the baseline T* approach. On the LV-Haystack benchmark, the average iteration count increases from 42.94 (T*) to 48.82 iterations, representing a modest 13.69% rise in the time cost.
Figure 4: Distribution of semantic-logical relations identified by VSLS in LongVideoBench and Video-MME datasets. Spatial relations dominate both (57.7% vs. 72.9%), with notable differences in causal relations (33.7% vs. 19.2%) and attribute relations (4.7% vs. 3.1%). Time relations remain least frequent (3.8% vs. 4.9%).
4.2 Ablation Study of Four Relations
Figure 4 illustrates the distribution of four logic relation types across LongVideoBench and Video-MME datasets, where spatial relations predominate, followed by attribute relations. In Table 4, we extract samples containing different relation types from LongVideoBench to compare the object detection-based T* method with our VSLS approach. Experimental results demonstrate that VSLS achieves significant improvements across both image similarity metrics (SSIM Precision and SSIM Recall). Additionally, temporal coverage shows marked enhancement for attribute, spatial, and causal relations, with spatial relations exhibiting the most substantial improvement (21.3% increase over T*). For the time relation category, we observe a slight decrease in temporal coverage, which may be attributed to the relative scarcity of time relation samples in the dataset, limiting the opportunity to demonstrate the advantages of VSLS. Nevertheless, Figure 1 provides visual evidence of how effectively leveraging time relations can facilitate downstream question-answering tasks.
LongVideoBench
Logic Type Method Precision $\uparrow$ Recall $\uparrow$ TC $\uparrow$
T* 72.9 88.7 37.5
Spatial VSLS (ours) 73.6 91.4 45.5
T* 71.8 87.6 38.5
Attribute VSLS (ours) 72.7 90.9 42.1
T* 76.7 89.2 37.3
Time VSLS (ours) 77.5 92.5 36.1
T* 74.7 92.4 38.6
Casual VSLS (ours) 74.7 93.8 39.6
Table 4: Results of searching utility on LongVideoBench. Precision: SSIM Precision; Recall: SSIM Recall; TC: Temporal Coverage. Results demonstrate significant improvements in two image-similarity metrics across all four categories of samples.
5 Related Work
Challenges in Long Video Understanding: Long video understanding is inherently more challenging than short-video or image-based tasks due to its rich temporal dynamics and massive redundancy [22, 43, 42]. The large number of frames increases both memory and computational requirements, making straightforward dense sampling infeasible. Moreover, crucial events may span distant timestamps, demanding high-capacity models to capture long-range dependencies [25, 26, 3, 34]. Meanwhile, the diverse and continuous visual content raises noise and distractors; thus, strategies to effectively locate or distill essential parts of the video are of primary importance [44, 7, 37, 39].
Existing Solutions based on VLMs typically share three core ideas: 1) video sampling or retrieval for efficiency, 2) multi-stage or interactive reasoning to handle complex questions, and 3) compact representation to accommodate the VLM’s limited context window. For instance, retrieval-based pipelines partition a video into segments and employ a learned or rule-based retriever to identify the relevant chunks before passing them to a VLM [20, 8, 9, 13]. Other lines of research compress each frame into minimal tokens to reduce computational overhead [16, 2, 27], or adopt a streaming mechanism to propagate memory representations along the temporal axis [22, 35, 19]. Beyond these efficiency-oriented approaches, LLM/VLM-as-planner frameworks factorize the process into a series of perception queries, enabling an agent to fetch additional frame-level details if needed [31, 45, 18].
6 Conclusion
In this paper, we present Visual Semantic-Logical Search (VSLS), a novel framework that efficiently selects semantically keyframes for long video understanding by decomposing logical relationships between textual queries and visual elements. VSLS based on four defined logical dependencies (spatial co-occurrence, temporal proximity, attribute dependency, and causal order), significantly outperforms existing methods while sampling only 1.4% of video frames. The 8.7% improvement in GPT-4o’s long video QA accuracy demonstrates that query-guided visual semantic logic search effectively bridges the gap between textual queries and visual content. VSLS’s plug-and-play nature enables seamless integration with existing pipelines, making it practical for real-world applications. Future work could consider more logical relations, learnable search methods, enhancing interpretability, and exploring more downstream tasks.
Supplementary Material
Part I Appendix

### v2

@@ 3.1 Benchmark Datasets
The proposed VSLS is systematically evaluated across four benchmark datasets: a) LongVideoBench lvhdataset for assessing long-context video-language comprehension capabilities; b) Video-MME videomme as the first comprehensive benchmark for multimodal video analytics; c) Haystack-LVBench, extended from LongVideoBench with human-annotated frame index answers; and d) Haystack-Ego4D, derived from Ego4D with similar annotations. While LongVideoBench and Video-MME measure performance enhancement in QA accuracy, Haystack-Ego4D and Haystack-LVBench quantitatively evaluate keyframe selection accuracy through recall and precision metrics. Further details of datasets are provided in Appendix LABEL:app:details.
@@ 3.2.1 Evaluation Metrics for Search Utility
Our assessment framework emphasizes both effectiveness and efficiency. For search effectiveness, we use three metrics to compare model-predicted keyframes with human annotations, considering both individual frames and full sets—addressing the possibility of multiple valid keyframe sets per query. For frame-level comparison, we evaluate the alignment between a predicted frame $f_{\text{pt}}$ and a human-annotated frame $f_{\text{gt}}$ from two perspectives:
@@
where $T_{\text{pt}}$ and $T_{\text{gt}}$ denote the sets of predicted and ground truth timestamps, respectively. Here, $|N_{\text{gt}}|$ is the number of ground truth frames, $t_{\text{gt}}^{i}$ and $t_{\text{pt}}^{j}$ are the $i$-th ground truth and $j$-th predicted timestamps, respectively. $\delta$ is the temporal similarity threshold defining the maximum allowed time deviation, and $\mathbb{I}[\cdot]$ is the indicator function, returning 1 if the condition holds and 0 otherwise.
Visual Similarity is measured by the Structural Similarity Index (SSIM) ssim, capturing structural detail, luminance, and contrast between $f_{\text{pt}}$ and $f_{\text{gt}}$. For set-to-set comparison, the key challenge is defining inter-set similarity. We adopt Precision $P$ and Recall $R$ as complementary metrics: Precision checks whether each predicted frame matches any reference frame, while Recall ensures that all reference frames are represented. Given the ground truth set $F_{\text{gt}}={f^{j}{\text{gt}}}^{n}{j=1}$ and the predicted set $F_{\text{pt}}={f^{i}{\text{pt}}}^{m}{i=1}$, we define the multimodal retrieval quality metrics as follows:
$P(F_{\text{pt}},F_{\text{gt}})$ $=\frac{1}{|F_{\text{pt}}|}\sum_{f^{i}_{\text{pt}}\in F_{\text{pt}}}\max_{f^{j}_{\text{gt}}\in F_{\text{gt}}}\phi(f^{i}_{\text{pt}},f^{j}_{\text{gt}}),$ (11a)
@@ 3.2.2 Evaluation Metrics for Search efficiency
Existing studies 13; 47; 65; 68; 71 have mainly concentrated on optimizing task-specific performance metrics while neglecting computational efficiency in temporal search operations. To systematically analyze this dimension, our evaluation framework incorporates two criteria: 1) FLOPs representing arithmetic operation complexity, and 2) Latency recording real-world execution duration.
@@ 3.3 Evaluation of Search Framework efficiency
[…]
[gray]0.7
[…]
Table 1: Evaluation of performance metrics across the LV-Haystack benchmark, presenting both search efficiency and end-to-end processing overhead (combining search and inference stages).
@@ 3.4 Visual Semantic Logic Search Performance
Table 2: Search utility results on LongVideoBench. Best scores in the 8-frame setting are underlined, and in the 32-frame setting are bold. Gray indicates results from the original paper.
[…]
These results empirically support our core hypothesis: leveraging semantic and logical cues from text queries enables precise detection of relevant video frames. Improvements in visual similarity and temporal coverage confirm that VSLS effectively captures keyframes while preserving temporal coherence through visual-logical alignment.